Shan lives 2.5 km from school. Jin lives 560 m from Shan's house. If the girls meet at Shan's house before walking to school, what distance does each girl walk
Answer by math_tutor2020(3817)


560 meters = 560/1000 = 0.56 km

Jin walks 0.56 km to get to Shan's house.
Then Jin walks another 2.5 km to get to school.
That's a total of 0.56+2.5 = 3.06 km

Meanwhile, Shan only walks 2.5 km since she of course starts at her own house.

Summary:
Jin = 3.06 km
Shan = 2.50 km